1. The force of +50.0μC at (-10.0m, 0) and +50.0μC at (10.0m, 0) on 20.0μC at (0, 0). Show solution.


2. The force of +40.0μC at (0, -3.00m) and +20.0μC at (0, 3.00m) on 50.0μC at (4.00m, 0). Show solution.


1
Expert's answer
2021-04-05T11:08:37-0400

"F=F_{13}-F_{23}=k{q_1q_3}\/{r_{13}^2}-kq_2q_3\/r_{23}^2="


"=9\cdot10^9\cdot50\cdot10^{-6}\cdot20\cdot10^{-6}/10^2-"


"-9\cdot10^{-9}\cdot50\cdot 10^{-6}\cdot20\cdot10^{-6}/10^2=0" . Answer




"F=\sqrt{F_{13}^2+F_{23}^2+2F_{13}F_{23}\cos73.74°}"



"F_{13}=k{q_1q_3}\/{r_{13}^2}="


"=9\cdot10^9\cdot40\cdot10^{-6}\cdot50\cdot10^{-6}/5^2=0.72\ (N)"



"F_{23}=kq_2q_3\/r_{23}^2="


"=9\cdot10^9\cdot20\cdot10^{-6}\cdot50\cdot10^{-6}/5^2=0.36\ (N)"



"F=\sqrt{0.72^2+0.36^2+2\cdot0.72\cdot 0.36\cdot\cos73.74°}="



"=0.89\ (N)" . Answer